Integrate the following functions w.r.t. x using substitution: 

 i) x x 1dx

1


3 4

+ ii) ∫

π

4


7 2

tan x sec x dx

 iii) ∫

sec (cos )x sin x dx


1
Expert's answer
2020-12-15T11:07:00-0500

1) Given "I=\intop _0^1x^3.(x^4+1)dx"

Let, "z=x^4+1"

Then,"dz\/dx =4x^3"

"\implies dz/4=x^3dx"

Now,"\intop x^3.(x^4+1)dx=\intop z.(dz/4)"

"=(1\/4).(z^2\/2)+C"

"=(1\/8).(x^4+1)^2+C"

Therefore, "I=\intop _0^1x^3.(x^4+1)dx"

"=[(1\/8).(x^4+1)^2+C]_0^1"

"=(1\/8).(1^4+1)^2-(1\/8).(0^4+1)^2"

"=4.(1\/8)-(1\/8)"

"=3\/8"


2)Given "I=\intop _{0}^{\pi /4} tan^7x.sec^2x dx"

Let, "z=tanx"

Then,"{dz}\/{dx}=sec^2x"

"\implies dz=sec^2xdx"

Again,when "x\to \pi /4,z\to1" and

when "x\to 0,z\to 0"

Therefore, "I=\intop _0^{\pi/4}tan^7x.sec^2xdx"

"=\intop _0^1z^7dz"

"=[z^8\/8]_0^1"

"=(1\/8).[1^8-0]"

"=(1\/8)"

3) Given "I=\intop sec(cosx).sinx" "dx"

Let, "z=cosx"

Then, "dz\/dx=(-sinx)"

"\implies (-dz)=sinx" "dx"

Therefore, "I=\intop sec(cosx).sinx" "dx"

"=\intop secz (-dz)"

"=-\intop secz" "dz"

"=-log|secz+tanz|+C"

"=-log|sec(cosx)+tan(cosx)|+C"